The length of a rectangular room is 3 feet longer than twice the width. If the room's perimeter is 210 feet, what are the room's dimensions?
Answer by redpeace(2)

Known: Perimeter=210 feet
Length= 2*n+3
Width= n
Equation = Perimeter= length + width
210= 2*n+3 + n
210= 2n+3+n
210=3n +3
210-3=3n
207=3n (divide both sides by 3)
207/3 = 3n/3
69=n
Length - 2*n +3 (2*69 +3) = 138+3 = 141 ft.
Width = 69 ft.
Perimeter - 141+69 = 210